5. A neutron star of mass 4.0 x 1030kg has a radius of about 12km. Calculate its mean density . comment on the answer.


1
Expert's answer
2021-02-17T11:07:43-0500
"\rho=\frac{m}{V}\\\rho=\frac{3m}{4\pi r^3}\\\rho=\frac{3(4\cdot10^{30})}{4\pi (12000)^3}=5.5\cdot10^{17}\frac{kg}{m^3}"